1. The person in Fig. 1 weighs 90 kg and the mass of the wagon (excluding the rocks) is 100 kg. Initially there are 50 one-kg rocks in the wagon. At t 0 the person starts to throw rocks out the back of the wagon at the rate of 1 rock per second. If the rocks leave the persons hand at a speed of 30 m/s, plot the position, X, and velocity U, of the wagon for t-1 to 50 seconds. Neglect wind drag on the person and the wagon. The initial conditions are: X 0 m and U 0 m/s at t-0 s (10 points) Case (a) Assume no friction loss.

Z,W x,u аз X,U One-Person Rocket. An astronaut throwing rocks out the back of a wagon is a smpl example of a rocket. The astronaut uses his muscles to accelerate the rocks in one di rection, leading to an equal but opposite force on the wagon that pushes it in the opposite direction. Note: if you try this experiment, friction on the wagons wheels could keep the propulsive force from actually moving it. (Adapted from Understanding Space: An Intro duction to Astronautics ISellers, 19941) Grayity Fig. 1
Answer #1

Body1 Body 2 Z,W V1 30 m/s V2 аз X,U tiAs there are no external forces acting (Assumed that friction losses are zero) so we can apply Principle of Conservation of Linear Momentum.

Mass of remaining rocks at time t= [Mass(at t=0sec - t] kg= (50-t) kg

M1= Mass of Body 1= (Mass of Wagon+Mass of person+ Mass of remaining Rocks at that time)

= (100+90+50-t) kg

=(240-t) kg

V1=velocity of Body 1

M2=Mass of body 2= mass of rock thrown away= 1kg

As per convention given in question, Velocity towards right is negative and that towards leftward is positive

V2=velocity of Body 2= (-30)i

sum Mi*Vi=0

M1*V1+ M2*V2=0

(240-t)*V1 + 1*(-30i)=0

(240-t)*V1 -30i=0

(240-t)*V1 = 30i

V1 = [30/ (240-t)] i m/s

for first second (t=1)

V1(t=1sec)= [30/ (240-1)] i m/s

=0.1255 m/s [towards left as thats positive i direction]

Similarly we can calculate for any second by substituting value of t for that moment,

V1(t=2sec)= [30/ (240-2)] i m/s

=0.1260 m/s

V1(t=3sec)= [30/ (240-3)] i m/s

=0.1265 m/s

.........

V1(t=47sec)= [30/ (240-47)] i m/s

=0.1554 m/s

V1(t=48sec)= [30/ (240-48)] i m/s

=0.1562 m/s

V1(t=49sec)= [30/ (240-49)] i m/s

=0.1570 m/s

V1(t=50sec)= [30/ (240-50)] i m/s

=0.1579 m/s

For Calculation of X:

We know   rac{mathrm{d} S}{mathrm{d} t}= V

int_{s1}^{s2}dS=int_{t}^{t+1}V*dt

As we know V is constant for any particular second

S2-S1= Vt *(t+1-t)= Vt*(1)= Vt

At start (t=0) S0=0

S1-S0=Vt=1

S1-0=0.1255 meters

S1=0.1255 meters

S2-S1=Vt=2

S2-0.1255=0.1260

S2=0.2515

similarly we can calculate all the St

0.1271 0 126S o1260 125S SC 0.5051 stope will be continuously imoeasing fo step O.2515 o.255 + (Sec